# Phase-space negativity as a computational resource for quantum kernel methods

### v1

@@ V.1 Output states of linear optical networks
where $W^{(\bm{s}}_{\rho_{\rm{in}}}(\bm{\alpha})$ is the ($\bm{s}$)-PQD of the input state $\rho_{\rm{in}}$. This relation shows that, given the ($\bm{s}$)-PQD of the input state, one can efficiently compute the ($\bm{s}$)-PQD of the data-encoding states $\rho(x)$, and use Algorithm 1 to estimate the kernel function.
In practice, however, any LON is lossy and hence cannot be described by a unitary transfer matrix. An interesting feature of our formalism is that it provides a practical way to take into account losses, as common sources of error, and check their effects on the negativity of the PQDs, which is directly related to the efficiency of our kernel estimation algorithms. In general, any lossy LON can be modeled as a quantum operation consisting of single-mode loss channels sandwiched between two lossless LONs [32]
@@
where $I-\bm{s}=\bm{\eta}(I-\bm{t})$ and $\bm{\eta}=\text{diag}(\eta_{1},\dots,\eta_{m})$ in the case of diagonal matrices of ordering parameters. This relation together with Eq. (26) enable us to investigate the effect of losses on the negativity and range of the PQDs of data-encoding states (27), and check when efficient estimation of the kernel function is possible.
To further illustrate our formalism let us assume that our initial state is product $\rho_{\rm{in}}=\bigotimes_{k=1}^{m}\rho_{k}$ and losses can be modeled in terms of single-mode loss channels at the input of lossless LONs. We then consider examples of single-photon states and cat states as the initial states. Under these assumptions Eq. (27) becomes $\rho(x)=U_{\text{LON}}(x)\bigotimes_{k=1}^{m}\tilde{\rho}_{k}U^{\dagger}_{\rm{LON}}(x)$ where $\tilde{\rho}_{k}=\Lambda_{\eta_{k}}(\rho_{k})$. Note that losses in state preparation can also be incorporated into these loss channels at the input of LONs. In this scenario, assuming that $\bm{s}=s\bm{I}$ in Eq. (26), the ($\bm{s}$)-PQD of the state $\rho(x)$ can be written as
@@
Here, $\mathcal{N}\big{(}W^{(s)}_{\tilde{\rho}_{k}}\big{)}=\int\text{d}^{2}\bm{\beta}_{k}|W^{(s)}_{\tilde{\rho}_{k}}(\beta_{k})|$. Notice that $\mathcal{N}\big{(}W^{(s\bm{I})}_{\rho(x)}\big{)}=\prod_{k=1}^{m}\mathcal{N}\big{(}W^{(s)}_{\tilde{\rho}_{k}}\big{)}$ since lossless LONs preserve the negativity and hence the non-classicality of quantum states [33]. We can also efficiently generate samples from the probability distribution (30) by first sampling the components of $N$ complex vectors $\bm{\beta}_{1},\dots,\bm{\beta}_{N}$ from individual probability distributions $|W^{(s)}_{\tilde{\rho}_{k}}(\beta_{k})|/\mathcal{N}\big{(}W^{(s)}_{\tilde{\rho}_{k}}\big{)}$, $k=1,\dots,N$, and then calculating $\bm{\alpha}_{j}=\bm{\beta}_{j}V^{\dagger}(x)$. Given these samples, we then obtain the corresponding values for the ($\bm{s}$)-PQD of $\rho(x^{\prime})$, $W^{(-s\bm{I})}_{\rho_{\rm{in}}}\!\big{(}\bm{\alpha}_{j}V(x^{\prime})\big{)}$, and through Eq. (18) calculate the sample mean $\frac{1}{N}\sum_{j=1}^{N}E(\bm{\alpha}_{j})$.
@@
V.1.1 Example 1: Single-photon states
@@
which is non-negative if $s\leq 1-2\eta_{k}$. Moreover, if $s>1-2\eta_{k}$, the negativity is given by
@@
To examine the bound (31), we should also consider the maximum value of $W^{(-s)}_{\tilde{\rho}_{k}}(\alpha)$. This function has two extreme values at $|\alpha_{0}|^{2}=0$ and $|\alpha_{1}|^{2}=(1+s)(4\eta_{k}-1-s)/(2\eta_{k})$, which are given by
@@ V.1.2 Example 2: Cat states
$W^{(s)}_{\ket{\text{cat}_{k}}}(\alpha)$ $=\frac{1}{\pi(1-s)(1+e^{-2|\gamma_{k}|^{2}})}\Bigg{(}\!\exp\!\bigg{[}{-2\frac{|\alpha+\gamma_{k}|^{2}}{1-s}}\bigg{]}$
$+\exp\!\bigg{[}{-2\frac{|\alpha-\gamma_{k}|^{2}}{1-s}}\bigg{]}$ (34)
$+2e^{-2|\gamma|^{2}}\Re\bigg{(}\!\exp\!\bigg{[}{-2\frac{(\alpha+\gamma_{k})(\alpha^{*}-\gamma_{k}^{*})}{1-s}}\bigg{]}\bigg{)}\!\Bigg{)},$
@@
$W^{(s)}_{\tilde{\rho}_{k}}(\alpha)=\frac{1}{\pi(1-s)(1+e^{-2|\gamma_{k}|^{2}})}\Bigg{(}\!\!\exp\!\bigg{[}{-2\frac{|\alpha+\sqrt{\eta_{k}}\gamma_{k}|^{2}}{1-s}}\bigg{]}$
$\quad+\exp\!\bigg{[}{-2\frac{|\alpha-\sqrt{\eta_{k}}\gamma_{k}|^{2}}{1-s}}\bigg{]}$ (35)
$\quad+2e^{-2|\gamma_{k}|^{2}}\Re\bigg{(}\!\!\exp\!\bigg{[}\!{-2\frac{\left(\alpha+\sqrt{\eta_{k}}\gamma_{k}\right)\!\left(\alpha^{*}-\sqrt{\eta_{k}}\gamma_{k}^{*}\right)}{1-s}}\bigg{]}\bigg{)}\!\!\Bigg{)}.$
@@ V.2 Gaussian states
$K(x,x^{\prime})\!=\!\frac{\exp\!\big{(}{-\frac{1}{2}(\bar{\bm{r}}-\bar{\bm{r}}^{\prime})(\Sigma(x)+\Sigma(x^{\prime}))^{-1}(\bar{\bm{r}}-\bar{\bm{r}}^{\prime})^{T}}\big{)}}{2^{-m}\sqrt{\det\left(\Sigma(x)+\Sigma(x^{\prime})\right)}}.$
@@
V.3 Partially measured Gaussian states
@@
where $\rho_{G}(x)=U(x)\rho_{\rm{in}}(x)U(x)^{\dagger}$ is an $(k+m)$-mode Gaussian state and $\Pi(x)=\bigotimes_{j=1}^{k}\Pi_{j}(x)$ is a tensor product of (possibly non-Gaussian) POVM elements. Recall that a quantum-efficient encoding refers to the fact that such states may be efficiently prepared using a quantum computer, a property which can be summarised here by $\mathrm{Tr}[(\Pi(x)\otimes\mathbb{I}_{m})\rho_{G}(x)]\geq 1/(\mathrm{poly}(m))$.
@@
For any classical data $x$, let $\rho(x)$ be a quantum state encoding over $m$ modes obtained by performing a measurement of the first $k$ modes of a $(k+m)$-mode Gaussian state $\rho_{G}(x)$, as in Eq. (37). Let $\tau(x)$ denote the non-classical depth of $\rho_{G}(x)$ (see Definition 1) and let $\tau(x,x^{\prime})=\max(\tau(x),\tau(x^{\prime}))\in[0,\frac{1}{2}]$. Then, assuming that the encoding is quantum-efficient, Algorithm 2 provides an estimate of the quantum kernel $K(x,x^{\prime})=\mathrm{Tr}[\rho(x)\rho(x^{\prime})]$ with additive precision $\epsilon$ and success probability $1-\delta$ in time
@@
A non-trivial consequence of Theorem 1 is that the efficiency of the classical simulation is independent of the non-Gaussianity of the measurements: even though these can inject a lot of negativity in the prepared state, this negativity does not affect the classical simulability through Algorithm 2, because the POVM elements only contribute to defining the estimators in Algorithm 2 and their non-Gaussianity does not substantially change the range of these estimators. In particular, even when making very non-Gaussian measurements (such as detecting many photons), classical simulation is always efficient as long as only a few modes $k=O(\log m)$ are measured.
What about when the number of measured modes is larger? There, Theorem 1 shows that classical simulation is still efficient when the Gaussian state being measured has small non-classicality, as quantified by the non-classical depth (see Definition 1). This notion of non-classicality is directly related to the amount of losses necessary to make the state fully classical [18], i.e., with non-negative $P$ function, and for a Gaussian state it is related to the minimum eigenvalue of its covariance matrix, see Eq. (36). This quantity can in turn be bounded by a function of the squeezing parameters and the symplectic eigenvalues encoding the impurity of the corresponding Gaussian state.
@@
VI Conclusion and outlook
We have introduced a framework based on phase-space quasi-probability distributions for the classical estimation of quantum kernel functions in machine learning. Our sufficient conditions, based on non-classical depth and on the negativity of quasi-probability distributions, identify these types of phase-space non-classicality measures as essential resources for achieving computational advantages in quantum machine learning with kernel methods. Our formalism can also be used to investigate the effect of errors and imperfections in quantum machine learning devices by examining their impact on the negativity of quasi-probability distributions.
@@ VI Conclusion and outlook
We have applied our results to a wide variety of examples, and have identified a subtle interplay between the quantum computational resources at hand: if no phase-space negativity is involved, and in the case of Gaussian measurements in particular, quantum kernels based on partially measured Gaussian states can be efficiently estimated classically using our Algorithm 1 (see Section V.2). When phase-space negativity is present in the measurements, then our Algorithm 2 still allows for efficient classical estimation of the corresponding quantum kernels as long as the number of measured modes is small enough or the non-classical depth of the Gaussian states involved is small enough (see Section V.3). In other terms, quantum computational advantage for quantum kernel estimation is only possible in this setting by combining Gaussian non-classical resources (squeezing) and non-Gaussian resources (phase-space negativity). A similar situation arises in the context of Gaussian boson sampling [40], where squeezing is a necessary ingredient together with non-Gaussianity for quantum computational advantage through sampling [41, 42].
Our results could be extended in a few directions. It would be interesting to apply the presented framework to the case of non-linear optical approaches such as the optical Ising machine [43]. Moreover, it is interesting to generalize the presented approach to the case of discrete-variable quasi-probability distributions using frame theory [44, 45].
Acknowledgements. U.C. acknowledges inspiring discussions with M. Walschaers, M. Frigerio, F. Arzani, J. Davis, M. Garnier, H. Thomas, P.E. Emeriau, S. Mehraban and D. Hangleiter. R.G acknowledges discussion with Chrsitoph Simon at the early stage of this project.
@@ Appendix A Computation of phase-space quasi-probability distributions
where $\mathcal{N}=\mathcal{N}\big{(}W^{(\bm{r}_{0})}_{\rho(x)}\big{)}\prod_{l=1}^{L}\mathcal{N}\big{(}T^{(\bm{r}_{l},-\bm{r}_{l-1})}_{u_{l}(x)}\big{)}\prod_{k=L+1}^{2L}\mathcal{N}\big{(}T^{(\bm{r}_{k},-\bm{r}_{k-1})}_{u_{k-L}(x^{\prime})}\big{)}$ is the total negativity, and $\vec{\bm{\gamma}}=(\bm{\gamma}_{0},\dots,\bm{\gamma}_{2L})$ is the vector of $(2L+1)m$ complex numbers. Viewing this expression as a Markov chain, by sampling form the distribution $\big{|}W^{(\bm{r}_{0})}_{\rho(x)}(\bm{\gamma}_{0})\big{|}/\mathcal{N}\big{(}W^{(\bm{r}_{0})}_{\rho(x)}\big{)}$ associated to the initial state that is known to be product, as well as other conditional probability distributions associated to the transition functions, one can generate $N$ random samples $\vec{\bm{\gamma}}_{1},\dots,\vec{\bm{\gamma}}_{N}$. Then, using the estimator
@@
Notice that, in general, this estimation algorithm is not as optimal as the estimation algorithm in terms of the ($\bm{s}$)-PQDs of the data-encoding states, see Eq. (10). Indeed, the negativity at the output of a quantum circuit is a lower bound of the product of the negativities of circuit elements because non-classical processes may reduce the effects of each other, e.g., a squeezing process and an anti-squeezing process can cancel each other.

### v2

@@ 5.1 Output states of linear optical networks
where $W^{(\bm{s})}_{\rho_{\rm{in}}}(\bm{\alpha})$ is the ($\bm{s}$)-PQD of the input state $\rho_{\rm{in}}$. This relation shows that, given the ($\bm{s}$)-PQD of the input state, one can efficiently compute the ($\bm{s}$)-PQD of the data-encoding states $\rho(x)$, and use Algorithm 1 to estimate the kernel function.
In practice, however, any LON is lossy and hence cannot be described by a unitary transfer matrix. An interesting feature of our formalism is that it provides a practical way to take losses into account, as common sources of error, and check their effects on the negativity of the PQDs, which is directly related to the efficiency of our kernel estimation algorithms. In general, any lossy LON can be modeled as a quantum operation consisting of single-mode loss channels sandwiched between two lossless LONs [36]
@@
where $I-\bm{s}=\bm{\eta}(I-\bm{t})$ and $\bm{\eta}=\text{diag}(\eta_{1},\dots,\eta_{m})$ in the case of diagonal matrices of ordering parameters. This relation, together with Eq. (25) enable us to investigate the effect of losses on the negative volume and range of the PQDs of data-encoding states (26), and check when efficient estimation of the kernel function is possible. Note that an alternative approach presented in [15] is to express the PQD of the state (26) in terms of the PQD of the input state and the Gaussian transition function associated with the lossy LON (see Eq.(39) of Appendix A). However, we emphasize that our classical estimation algorithms are inherently different from the classical sampling algorithms in [15], despite the similarities between the formalisms. For the sampling algorithms to work, all the PQDs must be nonnegative, while our estimation algorithms always work and can be efficient even in the presence of negativity.
To further illustrate our formalism, let us assume that our initial state is product $\rho_{\rm{in}}=\bigotimes_{k=1}^{m}\rho_{k}$ and losses can be modeled in terms of single-mode loss channels at the input of lossless LONs. We then consider examples of single-photon states and cat states as the initial states. Under these assumptions, Eq. (26) becomes $\rho(x)=U_{\text{LON}}(x)\bigotimes_{k=1}^{m}\tilde{\rho}_{k}U^{\dagger}_{\rm{LON}}(x)$ where $\tilde{\rho}_{k}=\Lambda_{\eta_{k}}(\rho_{k})$. Note that losses in state preparation can also be incorporated into these loss channels at the input of LONs. In this scenario, assuming that $\bm{s}=s\bm{I}$ in Eq. (25), the ($\bm{s}$)-PQD of the state $\rho(x)$ can be written as
@@
Here, $\mathcal{N}\big{(}W^{(s)}_{\tilde{\rho}_{k}}\big{)}=\int\text{d}^{2}\bm{\beta}_{k}|W^{(s)}_{\tilde{\rho}_{k}}(\beta_{k})|$. Notice that $\mathcal{N}\big{(}W^{(s\bm{I})}_{\rho(x)}\big{)}=\prod_{k=1}^{m}\mathcal{N}\big{(}W^{(s)}_{\tilde{\rho}_{k}}\big{)}$ since lossless LONs preserve the negative volume and hence the non-classicality of quantum states [22]. We can also efficiently generate samples from the probability distribution (29) by first sampling the components of $N$ complex vectors $\bm{\beta}_{1},\dots,\bm{\beta}_{N}$ from individual probability distributions $|W^{(s)}_{\tilde{\rho}_{k}}(\beta_{k})|/\mathcal{N}\big{(}W^{(s)}_{\tilde{\rho}_{k}}\big{)}$, $k=1,\dots,N$, and then calculating $\bm{\alpha}_{j}=\bm{\beta}_{j}V^{\dagger}(x)$. Given these samples, we then obtain the corresponding values for the ($\bm{s}$)-PQD of $\rho(x^{\prime})$, $W^{(-s\bm{I})}_{\rho_{\rm{in}}}\!\big{(}\bm{\alpha}_{j}V(x^{\prime})\big{)}$, and through Eq. (18) calculate the sample mean $\frac{1}{N}\sum_{j=1}^{N}E(\bm{\alpha}_{j})$.
@@
5.1.1 Example 1: Single-photon states
@@
which is non-negative if $s\leq 1-2\eta_{k}$. Moreover, if $s>1-2\eta_{k}$, the negative volume is given by
@@
To examine the bound (30), we should also consider the maximum value of $W^{(-s)}_{\tilde{\rho}_{k}}(\alpha)$. This function has two extremal values at $|\alpha_{0}|^{2}=0$ and $|\alpha_{1}|^{2}=(1+s)(4\eta_{k}-1-s)/(2\eta_{k})$, which are given by
@@ 5.1.2 Example 2: Cat states
$W^{(s)}_{\ket{\text{cat}_{k}}}(\alpha)$ $=\frac{1}{\pi(1-s)(1+e^{-2|\gamma_{k}|^{2}})}\Bigg{(}e^{-2\frac{|\alpha+\gamma_{k}|^{2}}{1-s}}$
$+e^{-2\frac{|\alpha-\gamma_{k}|^{2}}{1-s}}$ (33)
$+2e^{-2|\gamma|^{2}}\Re\bigg{(}e^{-2\frac{(\alpha+\gamma_{k})(\alpha^{*}-\gamma_{k}^{*})}{1-s}}\bigg{)}\!\Bigg{)},$
@@
$W^{(s)}_{\tilde{\rho}_{k}}(\alpha)=\frac{1}{\pi(1-s)(1+e^{-2|\gamma_{k}|^{2}})}\Bigg{(}e^{-2\frac{|\alpha+\sqrt{\eta_{k}}\gamma_{k}|^{2}}{1-s}}$
$\quad+e^{-2\frac{|\alpha-\sqrt{\eta_{k}}\gamma_{k}|^{2}}{1-s}}$ (34)
$\quad+2e^{-2|\gamma_{k}|^{2}}\Re\bigg{(}e^{-2\frac{\left(\alpha+\sqrt{\eta_{k}}\gamma_{k}\right)\!\left(\alpha^{*}-\sqrt{\eta_{k}}\gamma_{k}^{*}\right)}{1-s}}\bigg{)}\!\!\Bigg{)}.$
@@ 5.2 Gaussian states
$K(x,x^{\prime})\!=\!\frac{e^{-\frac{1}{2}(\bar{\bm{r}}-\bar{\bm{r}}^{\prime})(\Sigma(x)+\Sigma(x^{\prime}))^{-1}(\bar{\bm{r}}-\bar{\bm{r}}^{\prime})^{T}}}{2^{-m}\sqrt{\det\left(\Sigma(x)+\Sigma(x^{\prime})\right)}}.$
@@
5.3 Partially measured Gaussian states
@@
where $\rho_{G}(x)=U(x)\rho_{\rm{in}}(x)U(x)^{\dagger}$ is an $(k+m)$-mode Gaussian state and $\Pi(x)=\bigotimes_{j=1}^{k}\Pi_{j}(x)$ is a tensor product of (possibly non-Gaussian) POVM elements. Recall that a quantum-efficient encoding refers to the fact that such states may be efficiently prepared using a quantum computer, a property which can be summarized here by $\mathrm{Tr}[(\Pi(x)\otimes\mathbb{I}_{m})\rho_{G}(x)]\geq 1/(\mathrm{poly}(m))$.
@@
For any classical data $x$, let $\rho(x)$ be a quantum state encoding over $m$ modes obtained by performing a measurement of the first $k$ modes of a $(k+m)$-mode Gaussian state $\rho_{G}(x)$, as in Eq. (36). Let $\tau(x)$ denote the nonclassical depth of $\rho_{G}(x)$ (see Definition 1) and let $\tau(x,x^{\prime})=\max(\tau(x),\tau(x^{\prime}))\in[0,\frac{1}{2}]$. Then, assuming that the encoding is quantum-efficient, Algorithm 2 provides an estimate of the quantum kernel $K(x,x^{\prime})=\mathrm{Tr}[\rho(x)\rho(x^{\prime})]$ with additive precision $\epsilon$ and success probability $1-\delta$ in time
@@
A nontrivial consequence of Theorem 1 is that the efficiency of the classical simulation is independent of the non-Gaussianity of the measurements: even though these can inject a lot of negativity in the prepared state, as measured by the negative volume, this negativity does not affect the classical simulability through Algorithm 2, because the POVM elements only contribute to defining the estimators in Algorithm 2 and their non-Gaussianity does not substantially change the range of these estimators. In particular, even when making very non-Gaussian measurements (such as detecting many photons), classical simulation is always efficient as long as only a few modes $k=O(\log m)$ are measured.
What about when the number of measured modes is larger? There, Theorem 1 shows that classical simulation is still efficient when the Gaussian state being measured has small non-classicality, as quantified by the non-classical depth (see Definition 1). This notion of non-classicality is directly related to the amount of thermal noise necessary to make the state fully classical [23], i.e., with non-negative $P$ function, and for a Gaussian state it is related to the minimum eigenvalue of its covariance matrix, see Eq. (35). This quantity can in turn be bounded by a function of the squeezing parameters and the symplectic eigenvalues encoding the impurity of the corresponding Gaussian state.
@@
6 Conclusion and outlook
We have introduced a framework based on phase-space quasi-probability distributions for the classical estimation of quantum kernel functions in machine learning. Our sufficient conditions for efficient classical simulation are based on negative volume, non-classical depth, and excess range of quasi-probability distributions, and identify phase-space negativity as an essential resource for achieving computational advantages in quantum machine learning with kernel methods. Our formalism can also be used to investigate the effect of errors and imperfections in quantum machine learning devices by examining their impact on the negativity of quasi-probability distributions.
@@ 6 Conclusion and outlook
Moreover, we have identified a subtle interplay between the quantum computational resources at hand: if no phase-space negativity is involved, and in the case of Gaussian measurements in particular, quantum kernels based on partially measured Gaussian states can be efficiently estimated classically using our Algorithm 1 (see Section 5.2). When phase-space negativity is present in the measurements, then our Algorithm 2 still allows for efficient classical estimation of the corresponding quantum kernels as long as the number of measured modes is small enough or the non-classical depth of the Gaussian states involved is small enough (see Section 5.3). In other terms, quantum computational advantage for quantum kernel estimation is only possible in this setting by combining Gaussian non-classical resources (squeezing) and non-Gaussian resources (phase-space negative volume). A similar situation arises in the context of Gaussian boson sampling [43], where squeezing is a necessary ingredient together with non-Gaussianity for quantum computational advantage through sampling [44, 45].
Our results could be extended in a few directions. The sample complexities of our classical simulation algorithms are naturally expressed using non-classical measures relating to phase-space negativity; it would be interesting to relate these measures to other existing ones such as quadrature coherence scale [46], which provides an estimation of the distance to the set of classical states, or stellar rank [47], according to which a classification of bosonic kernels for quantum machine learning was recently derived [48]. Another direction could be to use our framework to analyze in more details how specific imperfections in implementations of quantum kernel methods, and the SWAP test in particular [49, 50, 51], could ease classical simulability. One could also apply the presented framework to the case of non-linear optical approaches such as the optical Ising machine [52] or Kerr-based kernels [53]. Moreover, it would be interesting to generalize the presented approach to the case of discrete-variable quasi-probability distributions using frame theory [54, 55].
Acknowledgements. U.C. acknowledges inspiring discussions with M. Walschaers, M. Frigerio, F. Arzani, J. Davis, M. Garnier, H. Thomas, P.E. Emeriau, S. Mehraban and D. Hangleiter. R.G. acknowledges discussion with C. Simon at the early stage of this project. U.C. acknowledges funding from the European Union’s Horizon Europe Framework Programme (EIC Pathfinder Challenge project Veriqub) under Grant Agreement No. 101114899.
Appendix
@@ Appendix A Computation of phase-space quasi-probability distributions
where $\mathcal{N}=\mathcal{N}\big{(}W^{(\bm{r}_{0})}_{\rho(x)}\big{)}\prod_{l=1}^{L}\mathcal{N}\big{(}T^{(\bm{r}_{l},-\bm{r}_{l-1})}_{u_{l}(x)}\big{)}\prod_{k=L+1}^{2L}\mathcal{N}\big{(}T^{(\bm{r}_{k},-\bm{r}_{k-1})}_{u_{k-L}(x^{\prime})}\big{)}$ is the total negative volume, and $\vec{\bm{\gamma}}=(\bm{\gamma}_{0},\dots,\bm{\gamma}_{2L})$ is the vector of $(2L+1)m$ complex numbers. Viewing this expression as a Markov chain, by sampling form the distribution $\big{|}W^{(\bm{r}_{0})}_{\rho(x)}(\bm{\gamma}_{0})\big{|}/\mathcal{N}\big{(}W^{(\bm{r}_{0})}_{\rho(x)}\big{)}$ associated to the initial state that is known to be product, as well as other conditional probability distributions associated to the transition functions, one can generate $N$ random samples $\vec{\bm{\gamma}}_{1},\dots,\vec{\bm{\gamma}}_{N}$. Then, using the estimator
@@
Notice that, in general, this estimation algorithm is not as optimal as the estimation algorithm in terms of the ($\bm{s}$)-PQDs of the data-encoding states, see Eq. (10). Indeed, the negative volume at the output of a quantum circuit is a lower bound of the product of the negative volumes of circuit elements because non-classical processes may reduce the effects of each other, e.g., a squeezing process and an anti-squeezing process can cancel each other.